# MMSearch-Plus: A Simple Yet Challenging Benchmark for Multimodal Browsing Agents

### v1

@@ 5.2 Observations
In addition to the result analysis from Section 4 that focuses on search modes and difficulties, we highlight three themes that relates to world knowledge, context window and visual clue grounding.
Closed-source MLLMs exhibit broad world knowledge, especially geographic.
When prompted with rare or niche events (e.g., images from small or sparsely documented locales), some closed-source models correctly recognize the event and provide relevant context, indicating substantial factual knowledge beyond mainstream datasets. By category, the geography tasks achieve the highest accuracies, consistent with findings from GeoBench (ccmdi, 2023).
Limitations remain in grounded reasoning.
Despite memory advantages, models often struggle with spatial grounding—e.g., producing correct bounding boxes for downstream search. Enabling agent-driven cropping frequently deteriorates performance on our benchmark. Consequently, we did not run a full-scale experiment with an image-cropping tool.
o3 shows strong memory retention.
Across several examples, o3 retains context over long interaction windows—successfully coordinating over 10+10{+} search rounds (Figure 5) and considering $10\times 5$ (top-5) $\approx 50{+}$ retrieved items. This persistence improves factual consistency and reduces redundant tool calls, leading to smoother task completion in information-seeking settings.
Qwen sometimes cannot call tools properly.
The bottom row of Figure 5 shows a long-tailed distribution of Qwen’s image-search calls—occasionally spiking to many calls within a single trajectory—while text-search calls remain low. On the right, incorrect trajectories exhibit increasing verbosity with more calls, whereas correct ones do not benefit from additional calls and sometimes even shorten, reflecting unstable tool-use dynamics. In practice, we also observe execution failures: out of $311$ tasks in MMSearch-Plus, Qwen produced invalid tool calls on $45$ tasks, for a total of $421$ invalid calls. These failures likely trigger cascaded retries (inflating call counts) without improving evidence quality, contributing to the observed variance and degraded accuracy.
Figure 6: Representative error case studies. Each panel corresponds to a high-frequency category in our taxonomy (Figure 7). Panels include the prompt, ground-truth event/answer, assigned error type, and a shortened rollout. (a) Key information not extracted: the summarizer omits required facts from retrieved pages. (b) Relevance not verified: a visually plausible page is retrieved, but the predicted winner corresponds to the wrong event (e.g., 2021 vs. 2024 snooker match). (c) Question misunderstood: the browsing agent misreads multimodal context and answers prematurely. (d) May require video understanding: the required datum exists only in narration from the source video, but our current search agent framework does not require video input.
@@ 5.3 Error Analysis
We categorize failure cases using a manually curated taxonomy and report the distribution over categories in Figure 7. Labels are assigned by the authors (single dominant label per erroneous prediction) for Gemini-2.5-Pro under the full-rollout setting.
@@
The two most prevalent failure modes are no relevant information found and hallucination, responsible for $51.1\%$ and $11.5\%$ of errors, respectively ($62.6\%$ combined). The former typically reflects a retrieval failure: noisy or overly specific queries miss on-topic pages. The latter reflects vision-driven misattribution: the model latches onto an incorrect event from internal knowledge and either does not invoke search or discounts contradictory evidence in the retrieved pages. Figure 6 provides examples of four additional high-frequency types.
@@
Figure 7: Distribution of error types. Manually annotated on Gemini-2.5-Pro (full rollout).
@@
Two recurring patterns emerge: (i) there is no single “golden path”—agents can succeed via varied orders of search/crop/verify; and (ii) stronger multimodal internal knowledge narrows the practical search space, enabling targeted verification over broad exploration. Accordingly, evaluation and training should reward families of effective strategies rather than a single script. While augmenting multimodal internal knowledge can reduce exploration overhead, end-to-end reliability still depends on robust verification and tighter cross-modal grounding.
@@ 6 Related Work
7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) reached only 15.1% without search and 36.0% with rollout, while a strong open-source model (Qwen-2.5-VL 72B) achieved 0.0% without search and 6.9% with rollout. These results indicate that current systems struggle to (i) reliably read and localize micro-signals (badges, micro-text, uniforms, layout geometry), (ii) decide when and how to crop and reuse visual evidence, and (iii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Limitations.
While diverse, our data inherits biases from public web sources and the domains we sampled; some tasks still under-represent non-English or low-visual-density pages. Our evaluation focuses on general-purpose agents with standard tool access; specialized systems with richer UI control or domain knowledge may behave differently. Finally, we primarily target image–text browsing; videos and dynamic interfaces remain underexplored.
Appendix A Comparison with Previous Benchmarks
Benchmark $M$ Search Multi‑hop (>>2) Thinking w/ Image Multi‑Image
Text Image Multi‑turn
InfoSeek
Frames
MMSearch
FactualVQA
ChineseSimpleVQA
BrowseComp
BrowseComp‑ZH
Ours
Appendix B Detailed Dataset Statistics
The distributions of primary categories and secondary categories across MMSearch-Plus are reported in Table 3 and 4, respectively.
@@ Appendix B Detailed Dataset Statistics
Table 4: Secondary categories.
Secondary Category Count %
technology 29 9.3%
Team Sports and Match Results 28 9.0%
Countries and Cities 20 6.4%
Television 18 5.8%
Paper and Publication Details 17 5.5%
Experimental Data and Findings 16 5.1%
Movies 16 5.1%
Addresses and Locales 15 4.8%
Transportation Infrastructure 13 4.2%
vlog 12 3.9%
Landmarks and Natural Features 10 3.2%
General Sports Events and Competitions 10 3.2%
Methodological Research 9 2.9%
Game Worlds and Characters 8 2.6%
Game Mechanics 8 2.6%
Esports and Competitive Gaming 8 2.6%
Live Performances 6 1.9%
Cultural and Historical Geography 6 1.9%
Individual Sports and Competitions 6 1.9%
Artists and Bands 5 1.6%
Short Films & Documentaries 4 1.3%
Technological Application 4 1.3%
Racket Sports 4 1.3%
Theoretical and Conceptual Studies 4 1.3%
Gaming Culture and Community 4 1.3%
Channel Characteristics 3 1.0%
Songs and Albums 3 1.0%
Esports and Non-traditional Sports 3 1.0%
Video Game Identification and Content 3 1.0%
Motor and Cycling Sports 3 1.0%
Music Production and Equipment 2 0.6%
Product Review & Launch 2 0.6%
Computing & AI 2 0.6%
Space Technology 1 0.3%
Musical Instruments 1 0.3%
Gadgets & Robotics 1 0.3%
Conventions & Events 1 0.3%
Personal Experience 1 0.3%
Film & TV Personalities 1 0.3%
Innovative Consumer Electronics 1 0.3%
Simulation & Modeling 1 0.3%
Wearable Technology 1 0.3%
Lifestyle & Tips 1 0.3%
Total 311 100.0%
Appendix C Results
The accuracies for o3 and Gemini-2.5-Pro across search modes are reported in Table 5.
Table 5: Results on MMSearch+ (accuracy %). Breakdown by search mode.
Model None One Image Full Rollout
Human
o3 (2025-04-16) 15.1 19.3 36.0
Gemini-2.5-Pro 10.6 16.4 23.8
Appendix D Rollout Trajectory Statistics
The statistics for the rollout trajectories of o3 under full rollout search mode are reported in Tables 6,7,891011 and 12.
Table 6: Quantiles per Conversation (Key Metrics)
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 2 3 7 16 20 20
Text queries total 5 11 24.5 49 63 83.8
Assistant words 14 21 36 56 67 119.1
Table 7: Quantiles per Conversation — Correct
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 1.0 2.0 5.0 8.9 12.0 15.9
Text search calls 1.0 2.0 5.0 8.9 11.5 15.8
Image search calls 0.0 0.0 0.0 1.0 1.0 1.0
Text queries total 5.0 9.5 19.0 31.0 41.9 61.7
Assistant characters 104.8 139.5 186.5 280.9 413.8 499.5
Assistant words 14.8 20.0 29.3 46.9 66.0 79.3
Tool result items 52.0 89.0 185.0 294.8 424.4 566.3
Tool result characters 18197.5 33739.5 79523.5 109980.6 158017.3 226533.7
Tool result words 2316.5 4543.5 9799.5 14974.8 21038.1 32243.8
Table 8: Quantiles per Conversation — Incorrect
Metric 25th Pct 50th Pct 75th Pct 90th Pct 95th Pct 99th Pct
Search calls total 2.0 3.0 9.0 20.0 20.0 20.0
Text search calls 2.0 3.0 9.0 19.0 20.0 20.0
Image search calls 0.0 0.0 1.0 1.0 1.0 2.0
Text queries total 6.0 13.0 31.0 53.6 67.4 87.0
Assistant chars 94.5 147.0 231.5 368.6 439.1 938.2
Assistant words 14.0 22.0 37.0 59.0 69.3 162.4
Tool result items 68.0 136.0 274.0 494.4 649.8 754.6
Tool result chars 27562.5 48914.0 88466.0 190252.8 256395.3 325728.4
Tool result words 3122.0 6263.0 11937.0 25224.8 34574.4 40940.8
Table 9: Means per Conversation by Correctness (selected metrics)
Metric All Correct Incorrect
Total search calls 5.4 3.7 6.4
Text search calls 5.2 3.5 6.1
Image search calls 0.2 0.1 0.3
Text queries total 19.3 14.2 22.2
Assistant words 27.5 25.1 28.9
Tool result items 179.7 135.1 204.9
Table 10: Top Tool Calls by Name
Tool Count
Text search 1613
Image search 75
Table 11: Top Tool Calls — Correct
Tool Count
Text search 397
Image search 16
Table 12: Top Tool Calls — Incorrect
Tool Count
Text search 1216
Image search 59
Appendix E Detailed Error Types Statistics
In Table 13, the manual annotations for a pilot study on the rollout trajectories of o3 on a previous version of our dataset of about 300 data points are provided.
Error Type Percentage
No relevant information found 51.1%
Hallucination 11.5%
Key information not extracted 8.4%
Relevance not verified 6.9%
Failed to understand the image 6.1%
Refused to answer 4.6%
Multiple answers 3.1%
Question misunderstood 3.1%
Search not invoked 2.3%
May require video understanding 1.5%
Summarization error 1.5%
Total 100.0%
Table 13: Error type distribution (minimal merge).

### v2

@@ 5.1 Observations
Figure 5: Stepwise tool distribution for Gemini, o3 and Qwen under Full Rollout+SoM, capped at Step $12$ for visualization. We use search round and step interchangeably, and an MLLM can make multiple search queries in one round (details in $\xi$ 3).
Closed-source models retained broad geographic knowledge.
They often recognized rare locales and supplied plausible context; geography tasks were comparatively strong. Without search tools, o3 achieved 31.2%, while human annotators with a standard browser only achieved 20.3%. On other categories like sports and Vlog, the human baseline usually outperformed the tested MLLMs. We name this phenomenon the different comparative advantages of humans and MLLMs.
o3 exhibited strong long-context coordination, while Gemini leveraged fewer search calls.
On the one hand, o3 sustained 10+10{+} search rounds and reasoned over $\sim 50{+}$ retrieved items without excessive redundancy, contributing to its rollout gains. On the other hand, Figure 5 (left) shows that Gemini always answered on or before Step 9. However, more tool calls did not correlate with trajectory correctness: Figure 14 shows that incorrect trajectories typically contained more search calls than correct ones.
Format correctness of closed-source models’ tool calls, and Qwen’s tool-use instability.
All the closed-source models consistently produced valid tool calls. For Qwen, we observed $421$ invalid image-search calls across $45$ of $311$ tasks, inflating retries without quality improvements.
More text search calls than image search calls, and zoom-in operations were not always followed by a sub-image search call.
In Figure 5, all models made more text-search calls than image-search calls. In addition, MLLMs sometimes zoomed in on the SoM-cropped image subregion only to see it clearly, without making a subsequent image search using the sub-image as the query.
@@ 5.2 Error Analysis
Figure 6: Distribution of error types. Annotated on Gemini-2.5-Pro (full rollout).
We categorize failure cases using a manually curated taxonomy and report the distribution over categories in Figure 6. Labels are assigned by the authors (single dominant label per erroneous prediction) for Gemini-2.5-Pro under the full-rollout setting. Examples of incorrect task trajectories are provided in Figure 7.
@@
The two most prevalent failure modes are no relevant information found and hallucination, responsible for $51.1\%$ and $11.5\%$ of errors, respectively ($62.6\%$ combined). The former typically reflects a retrieval failure: noisy or overly specific queries miss on-topic pages. The latter reflects vision-driven misattribution: the model latches onto an incorrect event from internal knowledge and either does not invoke search or discounts contradictory evidence in the retrieved pages.
@@
Two recurring patterns emerge: (i) there is no single “golden path”—agents can succeed via varied orders of search/crop/verify; and (ii) stronger multimodal internal knowledge narrows the practical search space, enabling targeted verification over broad exploration. While augmenting multimodal internal knowledge can reduce exploration overhead, end-to-end reliability still depends on robust verification and tighter cross-modal grounding.
Figure 7: Representative error case studies. Each row illustrates an example from a high-frequency error type (Figure 6). Panels show the prompt, GT answer, and a shortened rollout. From top to bottom: (a) Key info not extracted. (b) Relevance not verified. (c) Question misunderstood.
@@ 6 Related Work
Classical vision–language benchmarks probe compositional and commonsense reasoning in images and text (Johnson et al., 2017; Krishna et al., 2016; Zellers et al., 2019; Mathew et al., 2021), with recent suites extending to long videos and interleaved modalities (Wu et al., 2024; Li et al., 2024; Hu et al., 2025; Jeong et al., 2025). Concurrent work on web agents studies planning, browsing, and evidence gathering in open worlds (Krishna et al., 2024; Wei et al., 2025; Geng et al., 2025; Li et al., 2025a; Tao et al., 2025). Multimodal web search benchmarks begin to couple vision with browsing, but often emphasize entity matching over fine-grained visual reasoning (Jiang et al., 2024; Wu et al., 2025). Closest to our setting, MM-BrowseComp targets image/video evidence encountered during search and reveals substantial headroom even for frontier models (Li et al., 2025b). A complementary line promotes “thinking with images” via improved grounding and RL-style training to elicit image-first reasoning (Wu & Xie, 2024; Zheng et al., 2024; Fan et al., 2025; Zheng et al., 2025b; Su et al., 2025a; Xia et al., 2025). We differ by targeting sparse, spatial–temporal visual cues that must be transformed into targeted queries and multi-step plans, providing a focused stress test for multimodal browsing agents.
7 Conclusion
This work introduces MMSearch-Plus, a benchmark that restores the genuinely multimodal pressures of web browsing by requiring agents to extract weak, localized visual cues and propagate them through noisy retrieval to out-of-image facts. Our Spatial–Temporal Extrapolation curation procedure preserves the persistence and difficulty profile of text-only browsing while compelling fine-grained visual reasoning and provenance verification. Evaluating a range of open and closed MLLMs within a model-agnostic search agent framework reveals a substantial headroom: the strongest agent we tested (o3) reached only 15.1% without search and 36.0% with rollout, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieved 0.0% without search and 6.9% with rollout. These results indicate that current systems struggle to (i) reliably read and localize micro-signals (badges, micro-text, uniforms, layout geometry), (ii) decide when and how to crop and reuse visual evidence, and (iii) maintain verifiable chains of evidence across long-horizon tool use. We hope MMSearch-Plus serves as a rigorous stress test and a common yardstick for progress on multimodal browsing.
Appendix A Use of LLMs
The manuscript text was polished for fluency using Grammarly after the authors completed a full draft. The authors then manually reviewed all suggestions to ensure their appropriateness. In addition, ChatGPT was used to generate the task illustration in Figure 1 (upper left), based on the real image in Figure 8.
Appendix B Example from MMSearch-Plus
Appendix C Comparison with Previous Benchmarks
Benchmark/ Dataset $M$ Search
Multi-hop
(>>2)
Thinking
w/ Image
Multi-
Image
Text Image
Multi-
turn
InfoSeek (Chen, 2023)
FRAMES (Krishna et al., 2024)
MMSearch (Jiang et al., 2024)
FactualVQA(Wu et al., 2025)
ChineseSimpleVQA (Gu et al., 2025)
BrowseComp (Wei et al., 2025)
BrowseComp-ZH (Zhou et al., 2025)
BrowseComp-VL (Geng et al., 2025)
MM-BrowseComp (Li et al., 2025b)
Ours
Appendix D Related Work
We present a more comprehensive literature review in the Appendix.
@@ Appendix D Related Work
Appendix E Detailed Dataset Statistics
@@
The distributions of primary categories across MMSearch-Plus are reported in Table 3.
Appendix F Set-of-Mark Implementation Details
Figure 9 shows a list of task images overlaid with Set-of-Mark bounding box annotations from MMSearch-Plus.
In our context, set-of-mark for a task image consists of bounding boxes of key visual entities. Each box has a unique index. In the first user turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference. The framework exposes:
  • zoom_in(index): return a cropped sub-image for localized reasoning;
  • image_search(index): launch an image search seeded by the cropped region (compatible with the standard image-search pathway).
We include SoM because baseline MLLMs were unreliable at proposing visually precise crops from scratch. Providing accurate marks reduces the burden on detection/localization and lets models focus on reading, matching, and reasoning. While automatic mark generation (e.g., strong open-vocabulary detectors) is conceivable, robust coverage is challenging due to long-tail entities common in our benchmark; therefore MMSearch-Plus uses human-curated boxes.
Refer to caption Q: What is the title of this paper? A: MUCAR: Benchmarking Multilingual Cross-Modal Ambiguity Resolution for Multimodal Large Language Models
Refer to caption Q: What is the name of the previous level? A: Exercise Hall
Refer to caption Q: What is the title of this video? A: Who Can Identify as a Native American?
Refer to caption Q: In Paris Fashion Week, what brand did the person in the image represent? In which year did the event occur, and what is the event’s name? A: StreetStyle Winter 2025 Hermes
Figure 9: Examples of task images overlaid with Set-of-Mark bounding box annotations from MMSearch-Plus.
Appendix G Evaluation Details
Appendix H Agent Behavior Analysis
This appendix presents detailed visualizations of tool usage patterns across the three multimodal agents: Gemini, o3, and Qwen. All figures show episodes capped at 20 steps for readability.
H.1 Episode Raster Plots
Figure 10 shows the tool usage patterns for the 40 longest episodes from each agent. Each row represents one episode, sorted by length (longest at top). Columns represent sequential steps, with colors indicating tool types: text search, image search, zoom-in, and answer. Dark blue strips represent an MLLM did not fully utilize 20 rounds and stopped early.
Figure 10: Episode raster plots showing tool usage patterns for the 40 longest episodes per agent. Each row is an episode (sorted by length), each column is a step.
H.2 Tool Transition Matrices
Figure 11 shows the Markov transition probabilities between tool types.
[…]
H.3 Zoom-In Behavior Analysis
Figures 12 and 13 analyze zoom-in behavior patterns across agents. o3 shows the most aggressive zooming behavior, while Qwen and Gemini are more conservative.
Figure 12: Distribution of first zoom-in step positions. o3 tends to zoom earlier in episodes compared to Gemini and Qwen.
Figure 13: Distribution of zoom-in frequency per episode. o3 exhibits the highest zoom-in frequency, with some episodes containing 10+ zoom operations.
Appendix I Rollout Trajectory Statistics
Figure 14 shows the distribution of number of search calls for correct and incorrect task trajectories (left) and the relation of the number of words in a model’s trajectory (excluding the words for tool calls) vs. the number of search calls (right) under the Full Rollout setting.
Appendix J Prompts
System Prompt